# Reinforcement Twinning: from digital twins to model-based reinforcement learning

### v3

@@ 2.4 Novelties of the proposed approach
The RT framework proposed in this work is, in essence, MBRL with a white box (physics-based) model trained with methods from data assimilation. This model allows for real-time interaction and can be seen as a digital twin because it is trained, in real-time, to reproduce the performance of a specific system. Concerning the training, we use an ensemble adjoint-based approach to compute the gradient of the cost function driving the model derivation. This gives remarkable robustness to noise. Variants of the algorithm using Ensemble Kalman filtering are currently being explored.
However, contrary to traditional data assimilation, we assume that all states are observable and the underlying dynamic is deterministic. Moreover, since the main role of the model is to provide input-output relations in an actuated system, the approach is closer to system identification than traditional data assimilation. Finally, contrary to MPC, we assume that not all inputs to the system can be predicted; hence, the model cannot ‘look ahead’. However, the ensemble approach to handle the unpredictable inputs allows for online planning. Extensions to an MPC formalism are left to future work.
@@ 3 Definitions and General Formulation
    $\mathcal{J}_{a}(\bm{w}_{a,mb})=\mathbb{E}_{\sim S(\bm{z},\bm{w}_{a,mb}}\biggl{\{}\int^{T_{o}}_{0}\mathcal{L}_{a}\big{(}\tilde{\bm{s}}(t),\bm{s}(t;\bm{w}_{a,mb})\big{)}\;\mathrm{dt}\,\biggr{\}}\,,$ (3)
@@ 4.1 The Model-Free loop: (1)-(2)-(3)-(4)
We consider a classic actor-critic reinforcement learning formalism (Sutton and Barto 2018, Bhatnagar et al. 2009). Although more recent variants exist, we here focus on the classic Deep Deterministic Policy Gradient (DDPG) algorithm proposed by Lillicrap et al. (2019) with the prioritized experience replay strategy by Schaul et al. (2015). This algorithm offers a good compromise between performance and simplicity, and it has been proven successful in several flow control tasks (Bucci et al., 2019; Pino et al., 2023).
The DDPG trains two functions while interacting with the system. The first, referred to as actor, aims to approximate the policy function $\pi_{mf}:\check{\bm{e}}\rightarrow\check{\bm{a}}$, which provides the action to be taken at a certain state. This policy is deterministic, but we add noise generated by a random process to explore action space.
The second, referred to as critic, aims to approximate the state action value function $q:(\check{\bm{e}},\check{\bm{a}})\rightarrow Q$; this provides the expected reward for taking action $\check{\bm{a}}$ while the system is in the error state $\check{\bm{e}}$ and then following a certain policy. We use $Q$ to denote the (unknown) true Q value and $q(\check{\bm{e}},\check{\bm{a}};\bm{w}_{q})$ to denote the approximation from the $Q-$network. Referring to the previous section and Figure 1, the parameters $\bm{w}_{a,mf}$ denote the weights and biases of the actor-network while $\bm{w}_{q}$ denotes the weights and biases for the critic network.
These are identified through an online optimization process, which is based on stochastic, momentum-accelerated gradient-based optimization as in classic deep learning (see Goodfellow et al. (2016)). The gradient driving the optimization of the policy reads (see Silver et al. (2014)).
@@
The expectation is evaluated on an ensemble of trajectories or transitions and is readily available using back-propagation on the actor and the critic networks. Its accuracy strongly relies on the accuracy of the critic network. The recursive nature of the state value action function allows for measuring the performances of the critic from the Bellman equation (Sutton and Barto, 2018). For the true Q function this reads:
@@
The gradient of (7) with respect to $\bm{w}_{q}$ can be easily computed using back-propagation on the critic network. However, the training of the critic network is notoriously unstable (Mnih et al., 2013b, 2015). The two classic approaches to stabilize the learning consist in (1) using an under-relaxation in the update of the weights and (2) keeping track of a large number of transitions by using a replay buffer. The use of both methods is discussed in Section 4.3.
Finally, the sampling from the buffer is carried out in the form of batches to evaluate the gradient $d\mathcal{J}_{q}/d\bm{w}_{q}$. The sampling prioritizes transitions with the largest TD-error, $\delta_{k}$, (Schaul et al., 2016) building a non-zero probability distribution for each sample proportional to it $p_{i}=|\delta_{i}|+\mathrm{tol}$, indicating with tol a small offset. This yields to the sampling probability of the $i-$th transition, $P(i)$ to be expressed as:
@@ 4.2 The Model-Based Loop (1)- (2)- (3)-(5)-(6)
This function could be derived from first principles, as described in section 5 for the selected test cases, or using a general-purpose function approximator such as ANNs. The distinction is irrelevant to the method presented in this section, and we note in passing that using an ANN to approximate $f$ in combination with the training techniques described in this section leads to neural ODEs (Chen et al., 2018).
The problem of training the agent and the closure law is finding the optimal parameters $\bm{w}_{a}$ and $\bm{w}_{p}$, respectively, using continuously provided data. The training data consists of a target trajectory $\tilde{\bm{s}}(t)$ for the control problem and measurements of the states $\check{\bm{s}}(t)$ for the identification problem. The performances of the assimilating agent are measured according to (2) and (3), respectively (see Figure 1).
@@
To maximize the sample-efficiency training strategy, the optimization of these functions is carried out using gradient-based optimization along with efficient computation of the gradients $d\mathcal{J}_{p}/d\bm{w}_{p}$ and $d\mathcal{J}_{a}/d\bm{w}_{a}$ using the adjoint method (Errico, 1997; Bradley, 2019; Cao et al., 2003). This method bypasses the need for computing the sensitivities $d\bm{s}/d\bm{p}$ and $d\bm{s}/d{\bm{a}}$. Moreover, because of the similarity in the cost functions (2) and (3), the adjoint problems share key terms. Concerning the gradient driving the system identification, the adjoint-based evaluation gives
@@
Given a sample of the temporal evolution of the exogenous inputs $\bm{z}_{k}=\bm{z}(t_{k})$, with $k=[0,n_{t}-1]$, we build an analytic approximation of its large-scale component $\bm{d}_{k}=\bm{d}(t_{k})$ using support vector regression (Smola and Schölkopf, 2004). Let $\bm{d}_{k}[j]\in\mathbb{R}$ denote the j-th component of the exogenous input sampled at time step $k$. We seek to derive a continuous function $\tilde{\bm{d}}[j](t)$ such that $\tilde{\bm{d}}[j](t_{k})\approx\bm{d}_{k}[j]$. This is a regression problem that we tackle with support vector regression. Hence, the continuous function is written as
@@
The regression model generates $N_{z}$ samples of possible exogenous disturbances. These are modelled as a set of $n_{z}$ Gaussian processes with mean $\tilde{\bm{d}}[j](t;\bm{\alpha}[j])$ and covariance matrices $\bm{\Sigma}_{\varepsilon}[j]$. We sample these processes $N_{z}$ times to obtain the exogenous disturbances in the virtual trajectories. This procedure is shown in Fig. 2, for a simulated exogenous disturbance with $N_{z}=10$.
@@ 50       Policy switching and weight cloning based on the decision tree (Figure 3)
  • Step 2: Critic Update and model-free policy update (lines 14-24). An optimization with $n_{Q}$ iterations is carried out on the critic network, followed by an optimization with $n_{A}$ iterations for the policy. Both use batches of $n_{b}$ transitions sampled from $\mathcal{D}_{\mathrm{L}}$. Concerning the training of the critic network (lines 14-19), this loss function and its gradient are computed from (7). This step is essentially a training of the critic network in a supervised learning formalism. Similarly to what was proposed by Lillicrap et al. (2019) for the original DDPG, the update of the critic is modulated by an under-relaxation $\tau$ using a target network (line 17). At every $n_{Q}$ iteration, the optimized critic re-evaluates the TD errors of the sampled transitions in $\mathcal{D}_{\mathrm{L}}$ (line 18). Limiting the TD updates for the current batch of trajectories allows for a pseudo-random update of the most critical transitions, according to the critic. In addition, this approach is also computationally cheaper than re-evaluating all the stored transitions in the buffer, usually in the order $\mathcal{O}(10^{6})$. Successively, the critic is then used to train the policy in lines 20-24. In this second stage, the gradient with respect to the cost function is computed using eq (4). Note that we use the notation $\Delta\bm{w}$ to denote the optimization update for the weights $\bm{w}$, such that the updating reads $\bm{w}\leftarrow\bm{w}+\Delta\bm{w}$. Therefore, the update in a simple gradient descent over a cost function $J(\bm{w})$, for example, reads $\Delta\bm{w}=-\eta\;dJ(\bm{w})/d\bm{w}$. This makes the notation independent of the choice of the optimizer. Differently from the original DDPG algorithm in Lillicrap et al. (2019), the updates of the critic and the policy networks are carried out via two separate loops rather than in a single one. In our numerical experiments, we found this approach to lead to more stable model-free policy updates.
@@
  • Step 4: Optimize digital twin on the real actions (lines 27-36). The goal of this step is to find the closure parameters $\bm{w}_{p}$ such that the digital twin’s prediction matches with the $n_{S}$ trajectories of the real system, stored in $\mathcal{D}_{\mathrm{S}}$. Consequently, the loss function $\mathcal{J}_{p}$ in (2) is minimized using a gradient-based optimization loop. This consists of $n_{G}$ iterations and computes the updates $\Delta\bm{w}_{p}$ from the gradient evaluated along the $n_{S}$ trajectories available in $\mathcal{D}_{\mathrm{S}}$. Again, fixing the numerical time stepping to the one of the real interaction avoids the need for action interpolation, but might become problematic for stiff problems. For each trajectory (loop in lines 27-33), a total of $N_{z}$ possible realizations of the exogenous input are considered (sampled from the $\mathcal{N}^{(l)}_{z}$ processes built in the previous step), and the gradient is computed using the adjoint method (see eq. 10). The associated updates are then averaged (line 32), and the optimization step is carried out (line 34). Each optimization step requires a total of $n_{S}\times N_{Z}$ forward and backward virtual episodes.
@@
We close with a few general remarks on the algorithms.
@@
Tthe $n_{S}$ hyperparameter in Steps 4 and 5 defines how much the digital twin should consider past experiences. In this work, we set $n_{S}$=1, thus focusing on the last trajectory solely, while having a variable $N_{z}$ to promote robustness against randomness on the exogenous inputs and take into account possible sampling errors which may occur in real-world scenarios.
@@
The idle policy continues its training through different mechanisms. No special considerations are needed if the idle policy is the model-based policy: this continues to be trained on the digital twin regardless of whether it is active on the real system or not. On the other hand, if the idle policy is model-free, the learning occurs indirectly through two mechanisms that establish a “demonstration” from the model-based to the model-free: (1) the storage of “expert” transitions in the memory buffer $\mathcal{D}_{\mathrm{L}}$ and (2) the continuous training of the value function Q. The off-policy nature of the reinforcement learning algorithm here implemented allows the agent to learn from policies that are exploratory or simply suboptimal according to the current Q function. In this case, when the model-free policy is idle, and the Q function continues to be updated by observing the model-based policy in action on the real system, a sort of “demonstration” mechanism is established. The fact that the model-free could learn by the demonstration and eventually overtake the lead as a live policy is arguably the most significant contribution of this work.
@@ 5.1 Wind Turbine Control
We consider the problem of regulating the rotor’s angular speed $\omega_{r}$ and the extracted power $P_{e}$ of a variable speed variable pitch horizontal turbine of radius $R$, subject to time-varying wind velocity $u_{\infty}$ (see Laks et al. 2009; Pao and Johnson 2009).
@@
Figure 4: Sketch of the main parameters involved in wind turbine control (3(a)) and reference power curve (3(b)), $\tilde{P}$ (in red), and rotor speed $\tilde{\omega}_{r}$ (in blue).
Particularly critical is the transition between regions 2 and 3 (region 2.5 in Fig.4,) where gusts or lulls could result in conflicting requirements between torque and pitch controllers. This region is typically handled via set point smoothing, mitigating transients when switching from one control logic to another (Bianchi et al., 2007; Abbas et al., 2020).
In this test case, we combine a Maximum Power Point Tracking (MPPT) controller in below-rated conditions (Bhowmik and Spee, 1999; Howlader et al., 2010) with a regulation control in above-rated conditions. The controller can act on both the generator torque and the pitch angle, receiving as sole input the generator speed $\omega_{g}$, as typically done in multi-megawatt wind turbines (Jonkman et al., 2009). The current wind speed $u_{\infty}$ acts as exogenous input, which we assume can be provided by real-time measurements. We thus collect a sequence $u_{\infty}(t_{k})$ with $t_{k}=k\Delta t$. To avoid high-frequency excitation of the system, we filter the generator speed signal using an exponential smoothing (Zheng and Jin, 2022) with the recursive relations:
@@
We employ this filter for the generator rotational speed, $\omega_{g}$ and the measured wind speed, $u_{\infty}$. From the latter, it is possible to define the target rotational speed $\tilde{\omega}_{g}$ and extracted power $\tilde{P}$ from the reference curves in Figure 4. In Region 2, these two variables are linked by the operating conditions, while in Region 3, these are defined by mechanical constraints. In below-rated conditions, the power is maximized if the rotor operates at its optimal tip-speed-ratio $\lambda_{r,*}=\omega_{r}R/u_{\infty}$. Thus, one can define the reference generator speed relying on the measured wind speed as
@@
where $N_{g}$ is the gearbox ratio. This variable is constrained in the range $\omega_{g,min}\leq\tilde{\omega}_{g}\leq\omega_{g,\mathrm{rated}}$, taking into account the machine specifications. Therefore, given $\tilde{\omega}_{g}$ the target rotational velocity, we define the tracking errors as ${e}_{1}(t)=(\check{\omega}_{g}-\tilde{\omega}_{g})/\gamma_{1}$, and its integral over the past rotor revolution, ${e}_{2}(t)=(\int_{t_{i}-T_{r}}^{t_{i}}e(t^{\prime})\;\mathrm{dt}^{\prime})/\gamma_{2}$, where $\bm{\gamma}=[\gamma_{1},\gamma_{2}]$ ensure that both inputs lie in [-1, 1], and consider these as the inputs of the control logic. We then define the actuation vector as
@@
though any differentiable smooth step fits the purposes of bounding the action space while ensuring differentiability. This allows to limit the actuations rate of change according to the actuators’ specifications, $\dot{\tau}_{g,max}$ and $\dot{\beta}_{max})$. We further assume that the controller can be informed by measurements of the turbine’s rotational velocity $\check{\omega}$ and the generator torque $\check{\tau}_{g}$. It is thus possible to define a tracking error on the power as $e_{P}=\check{\tau}_{g}\cdot\omega(t;\bm{w}_{a})-\tilde{P}$ and use this signal to measure the controller’s performance. The reference power $\tilde{P}$ follows the $K\omega^{2}$-law (Bossanyi, 2000) in under-rated conditions, setting
$\tau_{g}(t)=K\omega_{g}(t)\quad\quad K=\frac{\pi\rho R^{5}C_{p,max}}{2\lambda_{r,*}^{3}N_{g}^{3}}\,,$ (21)
@@
Disregarding aeroelastic effects on both the blades and the tower, the simplest model of the rotor dynamics is provided by the angular momentum balance, which leads to a first-order system for the turbine’s generator angular speed $\omega_{g}$:
$\dot{\omega}_{g}=\frac{N_{g}}{J}\big{(}\tau_{a}-\tau_{g}N_{g})\,,$ (25)
@@
[…]
@@
[…]

### v4

@@ 2.5 Novelties of the Proposed Approach
The RT framework proposed in this work is a hybrid learning approach in which the model-based part is based on optimal control and adaptive modelling using methods from data assimilation. The first novelty within the landscape of model-based methods is the use of physics-based models with closure laws adapted in real-time: the derived model allows for real-time interaction and adapts to the system evolution, becoming a digital twin of the real system. Although the model is here trained using an ensemble adjoint-based approach to compute the gradient of the cost function driving the model derivation, variants of the algorithm using Ensemble Kalman filtering are possible and currently being explored. Departing from traditional data assimilation, we assume that all states are observable and the underlying dynamic is deterministic. Moreover, since the main role of the model is to provide input-output relations in an actuated system, the approach is closer to system identification than traditional data assimilation.
The second novelty within the land-scale of hybrid learning is in the strong interaction between the model-based and the model-free control agents, which alternate the role of disciple and master. A system similar to HME decides which policy is promoted to master and used on the real system. However, both are continuously trained in parallel, using the digital twin as a playground and sharing the collected expertise. We show that these agents can learn from each other via different mechanisms.
@@ 3 Definitions and General Formulation
    $\mathcal{J}_{a}(\bm{w}_{a,mb})=\mathbb{E}_{\sim S(\bm{z},\bm{w}_{a,mb})}\biggl{\{}\int^{T_{o}}_{0}\mathcal{L}_{a}\big{(}\tilde{\bm{s}}(t),\bm{s}(t;\bm{w}_{a,mb})\big{)}\;\mathrm{dt}\,\biggr{\}}\,,$ (3)
@@ 4.1 The Model-Free loop: (1)-(2)-(3)-(4)
We consider a classic actor-critic reinforcement learning formalism (Sutton and Barto 2018, Bhatnagar et al. 2009). Although more recent variants exist, we here focus on the classic Deep Deterministic Policy Gradient (DDPG) algorithm proposed by Lillicrap et al. [2019] with the prioritized experience replay strategy by Schaul et al. [2015]. This algorithm offers a good compromise between performance and simplicity, and it has been proven successful in several flow control tasks [Bucci et al., 2019, Pino et al., 2023].
The DDPG trains two functions while interacting with the system. The first, referred to as actor, aims to approximate the policy function $\pi_{mf}:\check{\bm{e}}\rightarrow\check{\bm{a}}$, which provides the action to be taken at a certain state. This policy is deterministic, but we add noise generated by a random process to explore the action space.
The second, referred to as critic, aims to approximate the state action value function $Q:(\check{\bm{e}},\check{\bm{a}})\rightarrow\mathbb{R}$; this provides the expected reward for taking action $\check{\bm{a}}$ while the system is in the error state $\check{\bm{e}}$ and then following a certain policy.
We use $Q$ to denote the (unknown) true $Q$ value and $q(\check{\bm{e}},\check{\bm{a}};\bm{w}_{q})$ to denote the approximation from the $Q-$network. In line with the previous section and Figure 1, the parameters $\bm{w}_{a,mf}$ denote the weights and biases of the actor-network while $\bm{w}_{q}$ denotes the weights and biases for the critic network. These are identified through an online optimization process, which is based on stochastic, momentum-accelerated gradient-based optimization as in classic deep learning (see Goodfellow et al. [2016]). The gradient driving the optimization of the policy reads (see Silver et al. [2014]):
@@
The expectation is evaluated on an ensemble of trajectories or transitions and is readily available using back-propagation on the actor and the critic networks. Its accuracy strongly relies on the accuracy of the critic network. The recursive nature of the state value action function allows for measuring the performances of the critic from the Bellman equation [Sutton and Barto, 2018]. For the true $Q$ function, this reads:
@@
The gradient of (7) with respect to $\bm{w}_{q}$ can be easily computed using back-propagation on the critic network. However, the training of the critic network is notoriously unstable [Mnih et al., 2013b, 2015]. The two classic approaches to stabilize the learning consist in (1) using an under-relaxation in the update of the weights and (2) keeping track of a large number of transitions by using a replay buffer. The implementation of both methods is discussed in Section 4.3.
Finally, the sampling from the buffer is carried out in the form of batches to evaluate the gradient $d\mathcal{J}_{q}/d\bm{w}_{q}$. The sampling prioritizes transitions with the largest TD-error, $\delta_{k}$, [Schaul et al., 2016] building a non-zero probability distribution for each sample proportional to it, i.e. $p_{i}=|\delta_{i}|+\mathrm{tol}$, with $\mathrm{tol}$ a small offset. This yields to the sampling probability of the $i-$th transition, $P(i)$ to be expressed as:
@@ 4.2 The Model-Based Loop (1)- (2)- (3)-(5)-(6)
This function could be derived from first principles, as described in section 5 for the selected test cases, or using a general-purpose function approximator such as ANNs. The distinction is irrelevant to the method presented in this section, and we note in passing that using an ANN to approximate $f$ in combination with the training techniques described in this section leads to neural ODEs [Chen et al., 2018].
The problem of training the agent and the closure law consists in finding the optimal parameters $\bm{w}_{a}$ and $\bm{w}_{p}$, respectively, using continuously provided data. The training data consists of a target trajectory $\tilde{\bm{s}}(t)$ for the control problem and measurements of the states $\check{\bm{s}}(t)$ for the identification problem. The performances of the assimilating agent are measured according to (2) and (3), respectively (see Figure 1).
@@
To maximize the sample-efficiency training strategy, the optimization of these functions is carried out using gradient-based optimization along with efficient computation of the gradients $d\mathcal{J}_{p}/d\bm{w}_{p}$ and $d\mathcal{J}_{a}/d\bm{w}_{a}$ using the adjoint method [Errico, 1997, Bradley, 2019, Cao et al., 2003]. This method bypasses the need for computing the sensitivities $d\bm{s}/d\bm{p}$ and $d\bm{s}/d{\bm{a}}$. Moreover, because of the similarity in the cost functions (2) and (3), the adjoint problems share key terms. Concerning the gradient driving the system identification, the adjoint-based evaluation gives
@@
Given a sample of the temporal evolution of the exogenous inputs $\bm{z}_{k}=\bm{z}(t_{k})$, with $k=[0,n_{t}-1]$, we build an analytic approximation of its large-scale component $\bm{d}_{k}=\bm{d}(t_{k})$ using Support Vector Regression (SVR, Smola and Schölkopf [2004]). Let $\bm{d}_{k}[j]\in\mathbb{R}$ denote the $j$-th component of the exogenous input sampled at time step $k$, with $j=[0,n_{z}-1]$. We seek to derive a continuous function $\tilde{\bm{d}}[j](t)$ such that $\tilde{\bm{d}}[j](t_{k})\approx\bm{d}_{k}[j]$. This is a regression problem that we tackle with support vector regression. Hence, the continuous function is written as
@@
The regression model generates $N_{z}$ samples of possible exogenous disturbances. These are modelled as a set of $n_{z}$ Gaussian processes with mean $\tilde{\bm{d}}[j](t;\bm{\alpha}[j])$ and covariance matrices $\bm{\Sigma}_{\varepsilon}[j]$. We sample these processes $N_{z}$ times to obtain the exogenous disturbances in the virtual trajectories. An example of SVR regression is shown in Figure 2 together with $N_{z}=10$ random samples of the modelled exogenous disturbance.
@@ 50       Policy switching and weight cloning based on the decision tree (Figure 3)
  • Step 2: Critic Update and model-free policy update (lines 14-24). An optimization with $n_{Q}$ iterations is carried out on the critic network, followed by an optimization with $n_{A}$ iterations for the policy. Both use batches of $n_{b}$ transitions sampled from $\mathcal{D}_{\mathrm{L}}$. Concerning the training of the critic network (lines 14-19), this loss function and its gradient are computed from (7). This step is essentially a training of the critic network in a supervised learning formalism. Similarly to what was proposed by Lillicrap et al. [2019] for the original DDPG, the update of the critic is modulated by an under-relaxation $\tau$ using a target network (line 17). At every $n_{Q}$ iteration, the optimized critic re-evaluates the TD errors of the sampled transitions in $\mathcal{D}_{\mathrm{L}}$ (line 18). Limiting the TD updates for the current batch of trajectories allows for a pseudo-random update of the most critical transitions, according to the critic. In addition, this approach is also computationally cheaper than re-evaluating all the stored transitions in the buffer, usually in the order $\mathcal{O}(10^{6})$. Successively, the critic is then used to train the policy in lines 20-24. In this second stage, the gradient with respect to the cost function is computed using eq (4). Note that we use the notation $\Delta\bm{w}$ to denote the optimization update for the weights $\bm{w}$, such that the updating reads $\bm{w}\leftarrow\bm{w}+\Delta\bm{w}$. Therefore, the update in a simple gradient descent over a cost function $J(\bm{w})$, for example, reads $\Delta\bm{w}=-\eta\;dJ(\bm{w})/d\bm{w}$. This makes the notation independent of the choice of the optimizer. Differently from the original DDPG algorithm in Lillicrap et al. [2019], the updates of the critic and the policy networks are carried out via two separate loops rather than in a single one. In our numerical experiments, we found this approach to lead to more stable model-free policy updates.
@@
  • Step 4: Optimize digital twin on the real actions (lines 27-36). The goal of this step is to find the closure parameters $\bm{w}_{p}$ such that the digital twin’s prediction matches with the $n_{S}$ trajectories of the real system, stored in $\mathcal{D}_{\mathrm{S}}$. Consequently, the loss function $\mathcal{J}_{p}$ in (2) is minimized using a gradient-based optimization loop. This consists of $n_{G}$ iterations and computes the updates $\Delta\bm{w}_{p}$ from the gradient evaluated along the $n_{S}$ trajectories available in $\mathcal{D}_{\mathrm{S}}$. Again, fixing the numerical time stepping to the one of the real interaction avoids the need for action interpolation but might become problematic for stiff problems. For each trajectory (loop in lines 27-33), a total of $N_{z}$ possible realizations of the exogenous input are considered (sampled from the $\mathcal{N}^{(l)}_{z}$ processes built in the previous step), and the gradient is computed using the adjoint method (see eq. 10). The associated updates are then averaged (line 32), and the optimization step is carried out (line 34). Each optimization step requires a total of $n_{S}\times N_{Z}$ forward and backward virtual episodes. It is worth highlighting that the ensemble averaging at this step would be essential in cases for which (1) the system is chaotic and highly sensitive to random disturbances (hence producing ”largely different” trajectories for ”similar” inputs) or (2) the policy becomes an exploratory/identification policy rather than a control one (hence the goal is not to keep a trajectory but explore as much as possible the state-action space). On the other hand, if the action is solely driven by the goal of trajectory tracking control and the system is inherently a noise-rejection system, the interest in computing many trajectories and then averaging the associated gradients is lost. The presented test cases fall within this second category; future work will consider cases under the first category.
@@
We close with a few general remarks on the proposed algorithm.
@@
The $n_{S}$ hyperparameter in Steps 4 and 5 defines how much the digital twin should consider past experiences. In this work, we set $n_{S}$=1, thus focusing on the last trajectory solely, while having a variable $N_{z}$ to promote robustness against randomness on the exogenous inputs and to take into account possible sampling errors which may occur in real-world scenarios.
@@
The idle policy continues its training through different mechanisms. No special considerations are needed if the idle policy is the model-based policy: this continues to be trained on the digital twin regardless of whether it is active on the real system or not. On the other hand, if the idle policy is model-free one, the learning occurs indirectly through two mechanisms that establish a “demonstration” from the model-based to the model-free: (1) the storage of “expert” transitions in the memory buffer $\mathcal{D}_{\mathrm{L}}$ and (2) the continuous training of the value function Q. The off-policy nature of the reinforcement learning algorithm here implemented allows the agent to learn from policies that are exploratory or simply suboptimal according to the current Q function. In this case, when the model-free policy is idle, and the Q function continues to be updated by observing the model-based policy in action on the real system– a sort of “demonstration” mechanism is established. The fact that the model-free could learn by the demonstration and eventually overtake the lead as a live policy is arguably the most significant contribution of this work.
@@ 5.1 Wind Turbine Control
We consider the problem of regulating the rotor’s angular speed $\omega_{r}$ and the extracted power $P_{e}$ of a variable speed variable pitch horizontal turbine of radius $R$, subject to time-varying wind velocity $u_{\infty}$ (see Laks et al. 2009, Pao and Johnson 2009).
@@
Figure 4: Sketch of the main parameters involved in wind turbine control (4(a)) and reference power curve (4(b)), $\tilde{P}$ (in red), and rotor speed $\tilde{\omega}_{r}$ (in blue).
Particularly critical is the transition between regions 2 and 3 (region 2.5 in Fig.4,) where gusts or lulls could result in conflicting requirements between torque and pitch controllers. This region is typically handled via set point smoothing, mitigating transients when switching from one control logic to another [Bianchi et al., 2007, Abbas et al., 2020].
In this test case, we combine a Maximum Power Point Tracking (MPPT) controller in below-rated conditions [Bhowmik and Spee, 1999, Howlader et al., 2010] with a regulation control in above-rated conditions. The controller can act on both the generator torque and the pitch angle, receiving as sole input the generator speed $\omega_{g}$, as typically done in multi-megawatt wind turbines [Jonkman et al., 2009]. The current wind speed $u_{\infty}$ acts as exogenous input, which we assume can be provided by real-time measurements. We thus collect a sequence $u_{\infty}(t_{k})$ with $t_{k}=k\Delta t$. To avoid high-frequency excitation of the system, we filter the generator speed signal using an exponential smoothing [Zheng and Jin, 2022] with the recursive relations:
@@
We employ this filter for the generator rotational speed, $\omega_{g}$ and the measured wind speed, $u_{\infty}$. From the latter, it is possible to define the target rotational speed $\tilde{\omega}_{g}$ and extracted power $\tilde{P}$ from the reference curves in Figure 4. In Region 2, these two variables are linked by the operating conditions, while in Region 3, these are defined by mechanical constraints. In the below-rated conditions, the power is maximized if the rotor operates at its optimal tip-speed-ratio $\lambda_{r,*}=\omega_{r}R/u_{\infty}$. Thus, one can define the reference generator speed relying on the measured wind speed as
@@
where $N_{g}$ is the gearbox ratio. This variable is constrained in the range $\omega_{g,min}\leq\tilde{\omega}_{g}\leq\omega_{g,\mathrm{rated}}$, taking into account the machine specifications. Therefore, given $\tilde{\omega}_{g}$ the target rotational velocity, we define the tracking errors as ${e}_{1}(t)=(\check{\omega}_{g}-\tilde{\omega}_{g})/\gamma_{1}$, and its integral over the past rotor revolution, ${e}_{2}(t)=(\int_{t_{i}-T_{r}}^{t_{i}}e_{1}(t^{\prime})\;\mathrm{dt}^{\prime})/\gamma_{2}$, where $\bm{\gamma}=[\gamma_{1},\gamma_{2}]$ ensure that both inputs lie in [-1, 1], and consider these as the inputs of the control logic. We then define the actuation vector as
@@
though any differentiable smooth step fits the purposes of bounding the action space while ensuring differentiability. This allows to limit the actuation rate of change according to the actuators’ specifications, $\dot{\tau}_{g,max}$ and $\dot{\beta}_{max})$. We further assume that the controller can be informed by measurements of the turbine’s rotational velocity $\check{\omega}$ and the generator torque $\check{\tau}_{g}$. It is thus possible to define a tracking error on the power as $e_{P}=\check{\tau}_{g}\cdot\omega(t;\bm{w}_{a})-\tilde{P}$ and use this signal to measure the controller’s performance. The reference power $\tilde{P}$ follows the $K\omega^{2}$-law [Bossanyi, 2000] in under-rated conditions, setting
$\tau_{g}(t)=K\omega_{g}(t)^{2}\quad\quad K=\frac{\pi\rho R^{5}C_{p,max}}{2\lambda_{r,*}^{3}N_{g}^{3}}\,,$ (21)
@@
Disregarding aeroelastic effects on both the blades and the tower, the simplest model of the rotor dynamics is provided by the angular momentum balance, which leads to a first-order system for the turbine’s rotor’s angular speed $\omega_{r}$:
$\dot{\omega}_{r}=\frac{1}{J}\big{(}\tau_{a}-\tau_{g}N_{g})\,,$ (25)
@@
We consider the NREL 5-MW reference offshore wind turbine [Jonkman et al., 2009], commonly used as a reference in the literature [De Cillis et al., 2022a, b, Coquelet et al., 2022]. This turbine has $R=63$ m, $J=115926$ kg m2 and $\omega_{r,\mathrm{rated}}=1.27$ rad s-1 for $u_{\infty}\geq 11.4$ m s-1. The gearbox ratio is $N_{g}=97$. The actuation bounds of the pitch angle and torque actuators are respectively $(\beta_{\mathrm{min}},\beta_{\mathrm{max}})=(0,90)$ deg and $(\tau_{\mathrm{min}},\tau_{\mathrm{max}})=(0,43093)$ Nm. Their maximum rate of change are $\dot{\beta}_{max}=8$ deg s-1 and $\dot{\tau}_{g,max}=15000$ N m s-1, respectively. From these, we define the cut-off frequency as the harmonic associated with the highest rate of change, obtaining $f_{c,\tau}=0.05$ Hz and $f_{c,\beta}=0.04$ Hz. We set the filter smoothing coefficient according to the slowest pitch actuator, finding $\hat{\alpha}=0.98$.
@@
We take $T_{o}=600$ s for the observation time and assume a sampling of $f_{s}=4$ Hz for all the simulated instrumentation. We simulate the incoming wind speed time series using the stochastic, turbulent wind simulator TurbSim [Jonkman, 2006] considering a wind velocity with average $u_{\infty}=11$ m s-1 and turbulence intensity TI$=16\%$, thus oscillating between below and above rated. An example of velocity time series is shown in Figure 5(a) together with the associated filtered signal.